OpenStudy (anonymous):

Rewrite with a rational exponet. 3rd square root of 18

OpenStudy (anonymous):

\[\sqrt[3]{18}\] ?

OpenStudy (anonymous):

\[\sqrt[3]{18}=\sqrt[3]{18^{1}}=18^{1/3}\]